OpenStudy (anonymous):

Help with the derivative of n^2/(1.1)^n?

OpenStudy (zehanz):

So you have \(f(n)=\dfrac{n^2}{(1.1)^n}\). -To get the derivative, you will need the Quotient Rule: \(\left(\dfrac{f}{g}\right)'=\dfrac{f'g-fg'}{g^2}\). -The derivative of a power function \(n^p\) is \(p\cdot n^{p-1}\). -The derivative of an exponential function \(a^n\) is \(\ln a \cdot a^n\). Armed with this information, you should be able to give it a try!